12. The many identical residents of Whoville love drinking Zlurp. Each resident has the following willingness to pay for the tasty refreshment: /3

First Bottle

$5

Fourth Bottle

$2

Second Bottle

$4

Fifth Bottle

$1

Third Bottle

$3

Further Bottles

$0


a. The cost of producing Zlurp is $1.50, and the competitive suppliers sell it at this price. (The supply curve is horizontal.) How many bottles will each Whovillian consume? What is each person’s consumer surplus? 

b. Producing Zlurp creates pollution. Each bottle has an external cost of $1. Taking this additional cost into account, what is total surplus per person in the allocation you described in part (a)? 

c. Cindy Lou Who, one of the residents of Whoville, decides on her own to reduce her consumption of Zlurp by one bottle. What happens to Cindy’s welfare (her consumer surplus minus the cost of pollution she experiences)? How does Cindy’s decision affect total surplus in Whoville?


1
Expert's answer
2021-10-31T18:23:29-0400

Let's build supply and demand graphs




a.Each resident will consume 4.5 bottles - this is the point of equilibrium.

The consumer's surplus (or gain) is the difference in demand volumes between the maximum price he is willing to pay for the product; Pmax, and the one he really pays Pe:

5--1,5=3,5

"The consumer's surplus=\frac{1}{2}(5-1.5)\times4.5=7.875"

b.

S=1+1.5=2,5

new equilibrium:




"The consumer's surplus=\frac{1}{2}(5-2.5)\times3.5=4.375"


c.I f Cindy Lou only consumes 1 bottle of Zlurp, her consumer surplus is $2.50 (= $5 −

$1.50 − $1). Cindy’s decision reduces consumer surplus in Whoville by $1.50.